Rational Equations. Solve the equation.
6/(x-1)+(2x)/(x-2)=2

Answer by Susmitha(21)

6/(x-1)+2x/(x-2)=2 (given)
6(x-2)+2x(x-1)/(x-1)(x-2)=2
6x-12+2x^2-2x=2(x-1)(x-2)
2x^2+4x-12=2[x^2-2x-x+2]
2x^2+4x-12=2x^2 -6x+4
4x-12=-6x+4
4x+6x=4+12
10x=16
x=16/10
x=1.6
here x^2 indicates x square or square of x
you can check the answer by substituting the obtained value of x in the given equation.